# The Case for ‘Thick Evaluations’ of Cultural Representation in AI

### v1

@@ Abstract.
Generative AI image models have been increasingly evaluated for their (in)ability to represent non-Western cultures. We argue that these evaluations operate through reductive ideals of representation, abstracted from how people define their own representation and neglecting the inherently interpretive and contextual nature of cultural representation. In contrast to these ‘thin’ evaluations, we introduce the idea of ‘thick evaluations:’ a more granular, situated, and discursive measurement framework for evaluating representations of social worlds in AI images, steeped in communities’ own understandings of representation. We develop this evaluation framework through workshops in South Asia, by studying the ‘thick’ ways in which people interpret and assign meaning to images of their own cultures. We introduce practices for thicker evaluations of representation that expand the understanding of representation underpinning AI evaluations and by co-constructing metrics with communities, bringing measurement in line with the experiences of communities on the ground.
Generative AI, Global South, cultural representation, qualitative evaluations, text-to-image
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Empirical studies in HCIccs: Applied computing Law, social and behavioral sciences
1. Introduction
@@
Generative AI (genAI) models, despite their popularity, have been shown to fail at inclusively representing different cultures in generated output, both images (Qadri et al., 2023a; Mim et al., 2024; Mack et al., 2024) and text (Wang et al., 2024b; Pawar et al., 2024; Singh et al., 2024; Myung et al., 2025; Khanuja et al., 2021), much like failures of representation of other types of AI systems (e.g., Katzman et al., 2023; Chien and Danks, 2024; Wang et al., 2024a; Kay et al., 2015; Shelby et al., 2023; Crawford, 2017). To address these gaps, prior work has sought to evaluate the cultural representations within AI generated output, but with few exceptions (Ghosh, 2024; Qadri et al., 2023b), mostly through quantified, metricized approaches to representation such as statistical similarities and benchmark-style scoring (Zhang et al., 2024; Li et al., 2024). However, the use of these methods presumes that representation is an objective construct with an empirical, definitive ground truth that outputs can be compared against (e.g., Kannen et al., 2024; Zhang et al., 2024) (for a critique of ground truth, see Muller et al., 2021). Given limitations of these computational methods, evaluation of representation is reduced to basic recognition or factual generation of artifacts. Even when human feedback on representation is sought, it is solicited through narrow, constrained, quantitative scales from anonymized crowdworkers who often do not have the lived experiences to evaluate nuances of cultural representation of other cultures.
@@
However, this approach to measuring representation is in contravention to decades of scholarship in the social sciences that emphasizes the subjective nature of representation, where judgments about representation in visual media are constructed in conversation with the viewer’s lived experiences and the broader context within which an image is viewed and published. (Hall, 1989, 1997). The evaluation categories many AI researchers use are also abstracted from the experiences of communities, constructed by AI researchers without engaging with communities to understand what axes of representation might be salient in their contexts. For instance, while many representational evaluations focus on skin tone, skin tone based racial categories might not be salient categories of differentiation in many cultures. Thus, while existing approaches to evaluating representation in AI/ML might be useful for evaluating the accuracy of visual depictions of the physical world, they may not allow us to engage with communities’ diverse desires for representations of their social worlds.
@@
We argue that effectively evaluating cultural representation in AI images requires ‘thick evaluations’—a more granular, situated, discursively constructed approach to measurement, steeped in communities’ own understanding of appropriate cultural representation. Our analysis draws on the ‘thick vs thin’ dichotomy introduced by philosopher Gilbert Ryle (Ryle, 1968), later taken on by Clifford Geertz (Geertz, 2008) and others (Ortner, 1995; Love, 2013; Jackson, 2013; Riles, 2000) to characterize descriptions of social worlds. Ryle’s classic example of a wink illustrates this distinction: a “thin description” merely describes the physical act of closing one eye by focusing on observable details, while a “thick description” unpacks the social meaning of the act and its significance as a signal to interlocutors. Seen through this framework, emerging methods for evaluating cultural representation in AI images involve ‘thin evaluations’; i.e., suited to evaluating the observable aspects of the physical world contained in AI-generated images, but not necessarily the social signals embedded in that physical visuals.
@@
To develop a ‘thick evaluation’ approach for cultural representation in AI images, we turned to the communities represented in images we seek to evaluate, recognizing their expertise and stakes in visual representations of their cultures. We conducted workshops with 37 participants in 3 South Asian countries to study the culturally situated ways people assign meaning to, interpret, and evaluate the representation of their cultures in AI-generated images. We find that 1) people evaluate representation of social worlds not just through a singular category such as accuracy, but through multi-dimensional, fine-grained axes; 2) people’s goals for evaluating cultural representation are situated in their social context, negotiated through dialogue with others and in response to broader societal discourse about their cultures; and 3) people deploy situated social knowledge and experiences with social worlds to evaluate varying social meanings of images.
@@
These findings demonstrate that thin evaluations alone cannot measure thick constructs like cultural representation and raise critical questions about the adequacy of existing evaluation paradigms for measuring cultural representation in AI images. As evaluation standards are being created for AI evaluations, this is a crucial moment to call for a thicker AI evaluation practice that reflexively interrogates epistemological underpinnings of AI evaluation practices, fundamentally rethinks whose expertise is included in the evaluation of AI systems, and opens up space for more interpretive qualitative methods. We provide pathways for such thickness by showing how AI researchers can interrogate the construct of cultural representation and the ways that that unobservable construct is operationalized in evaluation methods (cf. Jacobs and Wallach, 2021). We provide empirical data on varied categories of representation people might evaluate images for, which shows the need for co-constructing evaluation methods with members of communities whose culture is being represented to bring measurement in line with their experiences of AI images. We thus help AI practice move towards an ecosystem of ‘thick’ and ‘thin’ evaluations, encouraging congruence between the construct being evaluated and the evaluation methods used. As AI technology is entering into the space of cultural production, our work suggests the need to develop new, thicker forms of evaluations of cultural representation, to better reflect how people consume and interpret images of their cultures.
2. Related Work
2.1. Evaluating Cultural Representation in AI
@@
In order to understand the nature of AI models’ failures of representation (e.g., Qadri et al., 2023a; Mack et al., 2024), researchers have developed methods for evaluating representational failures (e.g., stereotyping) of AI systems (e.g., Katzman et al., 2023; Chien and Danks, 2024; Wang et al., 2024a; Kay et al., 2015; Shelby et al., 2023; Crawford, 2017; Harvey et al., 2024). Primarily, these evaluations have focused on language models (e.g., Blodgett et al., 2021; Hosseini et al., 2023; Abbasi et al., 2019; Katzman et al., 2023; Chien and Danks, 2024; Wang et al., 2024a), but they have also begun to include image models. For instance, emerging approaches for evaluating representation in images include developing quantified benchmark scores for goodness of representation (e.g., diversity (Kannen et al., 2024)), through methods like creating statistical similarity between generated images and reference images (Zhang et al., 2024), calculating correlations in keywords (Li et al., 2024), measuring model frequency of generating stereotypical and offensive images of nationality groups (Jha et al., 2024), calculating differences in frequency between the most and least common identities referenced in a set of model outputs (Lahoti et al., 2023). Human feedback has been sought via, for instance, using anonymized crowdworkers to score cultural biases in images (Basu et al., 2023). For representation in images, these measures often focus on visual demographic diversity characteristics—for instance, Cho et al. (2022) use automated skin tone and gender presentation classifiers to evaluate diversity of generated images.
@@
However, this line of work represents a technosolutionist, positivist (Crabtree, 2024) conception of representation, treating it as something objective, stable across time and contexts, and quantifiable. For instance, using skin tone to measure diversity presumes skin tone is a useful proxy for race and race is a universal category of differentiation, which may not be appropriate in all global cultures being represented. In contrast, researchers at the intersection of HCI and FAccT have conducted qualitative evaluations of representation in genAI image models, identifying failures of representation through more participatory and community-centered methods, with evaluators who are from the same culture as the images they are evaluating (Qadri et al., 2023b; Mim et al., 2024; Ghosh et al., 2024). However, while these efforts have demonstrated the value of qualitative approaches for evaluating representation in generative image models, they do not examine what is meant by representation as a construct nor do they critically interrogate what the process of interpretive meaning-making that participants in such studies undertake suggests for broader approaches to evaluations of cultural representation in AI.
@@
Recent research has shed new light on the complex process of measuring unobservable social constructs such as fairness, as well as the potential harms of reductive measurement approaches. For instance, Selbst et al. (2019) critique the definitions of fairness that lead to misleading abstractions in its evaluation. Recognizing the contested nature of the construct of fairness, Smith et al. (2024) argue for bringing in stakeholders’ expertise for more meaningful and contextually grounded evaluations. Scholars like Jacobs and Wallach (2021) advocate for utilizing approaches to measurement modeling from the social sciences to bridge the gap between unobservable constructs and their operationalization in evaluations of fairness in algorithmic systems—a framework we draw on in this paper through our discussion of the construct of cultural representation and how it is operationalized in evaluations.
@@
Jacobs and Wallach (2021) argue that measurement modeling begins with a clear and robust articulation of the construct being evaluated—in their paper, fairness, and in our paper, cultural representation. However, in prior work evaluating cultural representation in text-to-image models (e.g., Zhang et al., 2024; Kannen et al., 2024; Jha et al., 2024; Lahoti et al., 2023; Basu et al., 2023), the construct of representation is often not interrogated or explicitly defined. Although often left implicit, prior work operationalizes representation in various ways, as lack of bias (Wan et al., 2024), as geo-cultural similarity (Basu et al., 2023), and as imbalance (Su et al., 2009). More broadly than for cultural representation in text-to-image models, Chasalow and Levy (2021) argue that researchers often fail to make explicit how they conceptualize ‘representative’ or ‘representation.’ Such various ways of operationalizing the same construct may lead to evaluations that optimize for disparate goals (e.g., cultural awareness, cultural diversity, avoiding “cultural bias” (Li et al., 2024), etc), making it difficult to compare the effectiveness of different approaches (Jacobs and Wallach, 2021).
2.2. Situated Conceptualizations of Representation
@@
Decades of scholarship from the humanities have argued that representation in media, particularly visual media, cannot be defined through positivist approaches (e.g., Hall, 1989, 1997; Barthes, 1999). They argue that representation is not a static goal that can be objectively and computationally evaluated and achieved, but is instead an ongoing interpretive act whose meaning is contested and negotiated. In media studies, visual media, such as paintings, photography, or digital art, do not represent some objective reality but instead communicate multiple shades of meaning about the world, known as polysemy (Hall, 1997). That is, images have both denotative meanings (i.e., the literal subject of the image) as well as connotative meanings (i.e., the symbolic meaning) the image was designed to convey—or the meaning it invokes in the viewers, which may be different than the intended meaning of the image (Hall, 1997; Barthes, 1999).
@@
As a result, for media studies scholars, representation in visual media is a process of meaning-making, not a one-to-one depiction of the world as it is—and thus representation is a site for struggle over meaning (Hall, 1989, 1997; Desai, 2000). In this way, even a photograph of an event does not objectively convey a singular true meaning (Hall, 1997), but instead conveys a “positional truth” (Abu-Lughod, 1991) (or multiple such truths, given the polysemic nature of images), shaped by the creators’ and viewers’ cultural and historical context. Thus, visual understanding itself relies on a slippery relationship between words, images, and the concepts signified by those images (Barthes, 1999). Indeed, as media studies scholars remind us, representation in public culture is a “zone of contestation” (Desai, 2000), characterized by contests over which peoples and perspectives are made visible, how they are portrayed (and by whom), contests which are shaped by who has the power to control narratives of representation. Thus, in this paper, we explore how people evaluate representation in AI-generated images and what this representation means to them in their culture.
2.3. Thick vs Thin Modes of Understanding the Social
The ‘thick/thin’ dichotomy, originating in Gilbert Ryle’s philosophical work (Ryle, 1968), has become a cornerstone of anthropological and social scientific inquiry. He argued that understanding human behavior requires moving beyond mere observation of physical details to interpretation, recognizing the “many-layered sandwich” of meaning embedded within even seemingly simple actions (Ryle, 1968). This concept was further elaborated upon by Clifford Geertz, who emphasized the importance of thick description in ethnographic research (Geertz, 2008). Geertz argued that researchers must not only describe actions but also interpret them within their social or cultural context. A wink, for instance, could signify a conspiratorial gesture, a flirt, a parody, or a mere rehearsal, each carrying distinct meanings depending on the social context.
@@ 2.3. Thick vs Thin Modes of Understanding the Social
Sherry Ortner, in a chapter on “thick resistance,” highlighted the limitations of thin approaches in capturing the complexities of social life, which require the “richness, texture, and detail,” of thickness (Ortner, 1995, 1997). She argued that thin descriptions often overlook the “internal politics of dominated groups, thin on the cultural richness of those groups, thin on the subjectivity—the intentions, desires, fears, projects—of the actors engaged in these dramas” (Ortner, 1995). Ortner advocated for thick descriptions that delve into the “production, circulation, and consumption” of cultural practices, recognizing their dynamic interplay with social and political realities. While the value of thickness in capturing the nuances of social phenomena has been widely acknowledged, scholars have also recognized the importance of thinness in discerning patterns and broader social structures. Annelise Riles (Riles, 2000), in her work on multi-sited ethnography, explored the challenges of achieving thick description when studying global phenomena dispersed across diverse cultures, calling attention to “the limits of thickness as a disciplinary trope” and to pay more attention to what the use of “thin composition” could yield. Anthropologist John Jackson similarly emphasized the value of thin description in providing a “baseline empiricism” and a starting point for social investigation (Jackson, 2013). Thus, while thick and thin both have their place in social analysis, thinness can only get you so far. In this paper, we draw on the thick/thin dichotomy to provide a framework for navigating the complexities of social inquiry when considering multiple approaches to evaluating cultural representation in AI images.
3. Methods

### v2

# The Case for “Thick Evaluations” of Cultural Representation in AI

@@ Abstract
Generative AI model outputs have been increasingly evaluated for their (in)ability to represent non-Western cultures. We argue that these evaluations often operate through reductive ideals of representation, abstracted from how people define their own representation and neglecting the inherently interpretive and contextual nature of cultural representation. In contrast to these ‘thin’ evaluations, we introduce the idea of ‘thick evaluations’: a more granular, situated, and discursive measurement framework for evaluating representations of social worlds in AI outputs, steeped in communities’ own understandings of representation. We develop this evaluation framework through workshops in South Asia, by studying the ‘thick’ ways in which people interpret and assign meaning to AI-generated images of their own cultures. We introduce practices for thicker evaluations of representation that expand the understanding of representation underpinning AI evaluations and by co-constructing metrics with communities, bringing measurement in line with the experiences of communities on the ground.
Introduction
Generative AI (genAI) models, despite their popularity, have been shown to fail at inclusively representing different cultures in generated outputs—including images (Qadri et al. 2023; Mim et al. 2024; Mack et al. 2024) and text (Wang et al. 2024b; Pawar et al. 2024; Singh et al. 2024; Myung et al. 2025; Khanuja et al. 2021)—much like failures of representation in other AI systems (e.g., Katzman et al. 2023; Chien and Danks 2024; Wang et al. 2024a; Kay et al. 2015; Shelby et al. 2023; Crawford 2017).
@@ Introduction
To address these gaps, prior work has sought to evaluate the cultural representations in AI-generated output, but with few exceptions (e.g., Ghosh 2024; Qadri et al. 2023), mostly through quantified, metricized approaches to representation such as statistical similarities and benchmark-style scoring (Zhang et al. 2024; Li et al. 2024). The use of these methods presumes that representation is an objective construct with a quantifiable, definitive ground truth that outputs can be compared against (e.g., Kannen et al. 2024; Zhang et al. 2024) (for a critique of ground truth, see Muller et al. 2021). Given limitations of computational methods, representation is reduced to evaluations of basic recognition or factual generation of artifacts. When human feedback on representation is sought, it is often solicited through narrow, constrained, quantitative scales from crowdworkers who may not have the lived experiences to evaluate nuances of content’s varied meanings to different social groups (Díaz et al. 2022).
@@
However, this approach to measuring representation is in contravention to decades of scholarship in the social sciences that emphasizes the subjective nature of representation, where judgments about representation in visual media are constructed in conversation with the viewer’s lived experiences and the broader context within which an image is viewed and published. (Hall 1989, 1997). In addition, the categories many AI researchers use for evaluations are often abstracted from the experiences of communities, constructed by AI researchers without engaging with communities to understand what axes of representation might be salient in their contexts. For instance, while many representational evaluations focus on skin tone, skin tone might not be a salient category of differentiation for social groups in many cultures. Thus, while existing approaches to evaluating representation in AI/ML might be useful for evaluating the accuracy of visual depictions of the physical world, they may not allow us to engage with communities’ diverse desires for representations of their social worlds.
@@
We argue that effectively evaluating cultural representation in AI images requires ‘thick evaluations’—a more granular, situated, discursively constructed approach to measurement, steeped in communities’ own understanding of appropriate cultural representation. Our analysis draws on the ‘thick vs thin’ dichotomy introduced by philosopher Gilbert Ryle (1968), later taken on by Clifford Geertz (2008) and others (Ortner 1995; Love 2013; Jackson 2013; Riles 2000; Nelson 2023a, b) to characterize descriptions of social worlds. Ryle’s classic example of a wink illustrates this distinction: a ‘thin description’ merely describes the physical act of closing one eye by focusing on observable details, while a ‘thick description’ unpacks the social meaning of the act and its significance as a signal to interlocutors. Seen through this framework, emerging methods for evaluating cultural representation in AI images involve ‘thin evaluations’; i.e., suited to evaluating the observable aspects of the physical world contained in AI-generated images, but not necessarily the social signals embedded in those physical depictions.
@@
To develop a ‘thick evaluation’ approach for cultural representation in AI images, we turned to the communities represented in images we seek to evaluate, recognizing their expertise and stakes in visual representations of their cultures. Given prior work on AI’s failures of representation for South Asian cultures (e.g., Qadri et al. 2023; Ghosh et al. 2024b), we conducted workshops with 37 participants in 3 South Asian countries to study the culturally-situated ways people assign meaning to, interpret, and evaluate the representation of their cultures in AI-generated images. We find that 1) people evaluate representation of social worlds not just through a singular category such as accuracy, but through multi-dimensional, fine-grained axes; 2) people’s goals for evaluating cultural representation are situated in their social context, negotiated through dialogue with others and in response to broader societal discourse about their cultures; and 3) people use situated social knowledge of and experiences with social worlds to evaluate varying social meanings of images.
@@
These findings demonstrate that thin evaluations alone cannot measure thick constructs like cultural representation, and they raise critical questions about the adequacy of existing evaluation paradigms for measuring cultural representation in AI images. As new methods and standards are being created for AI evaluations, this is a crucial moment to call for a thicker AI evaluation practice that reflexively interrogates epistemological underpinnings of AI evaluation practices, fundamentally rethinks whose expertise is included in the evaluation of AI systems, and opens up space for more interpretive qualitative evaluation methods. We provide pathways for such thickness by showing how AI researchers can interrogate the construct of cultural representation and the ways that that unobservable construct is operationalized in evaluation methods (cf. Jacobs and Wallach 2021). We provide empirical data on varied categories of representation people might evaluate images for, which shows the need for co-constructing evaluation methods with members of communities whose culture is being represented, to bring measurement in line with their experiences of AI images. We thus help AI practice move towards an ecosystem of both ‘thick’ and ‘thin’ evaluations, encouraging congruence between the construct being evaluated and the evaluation methods used. As AI technology is entering into the space of cultural production, our work suggests the need to develop new, thicker forms of evaluations of cultural representation, to better reflect how people consume and interpret images of their cultures.
Related Work
Evaluating Cultural Representation in AI
@@
In order to understand the nature of AI models’ failures of representation (e.g., Qadri et al. 2023; Mack et al. 2024), researchers have developed methods for evaluating representational failures (e.g., stereotyping) of AI systems (e.g., Katzman et al. 2023; Chien and Danks 2024; Wang et al. 2024a; Kay et al. 2015; Shelby et al. 2023; Crawford 2017; Harvey et al. 2024). Primarily, these evaluations have focused on language models (e.g., Blodgett et al. 2021; Hosseini et al. 2023; Abbasi et al. 2019; Katzman et al. 2023; Chien and Danks 2024; Wang et al. 2024a; Wolfe et al. 2024), but they have also begun to include image models. For instance, emerging approaches for evaluating representation in images include developing quantified benchmark scores for goodness of representation (e.g., diversity (Kannen et al. 2024)), via methods like creating statistical similarity between generated images and reference images (Zhang et al. 2024; Ghosh 2024), calculating correlations in keywords (Li et al. 2024), measuring model frequency of generating stereotypical and offensive images of nationality groups (Jha et al. 2024), and calculating differences in frequency between the most and least common identities referenced in a set of model outputs (Lahoti et al. 2023). Human feedback has been sought via, for instance, using anonymized crowdworkers to score cultural biases in images (Basu et al. 2023; Ghosh et al. 2024a). For representation in images, these measures often focus on visual demographic diversity characteristics—for instance, Cho et al. (2022) use automated skin tone and gender presentation classifiers to evaluate diversity of generated images.
@@
However, this line of work follows a technosolutionist, positivist (Crabtree 2024) conception of representation; i.e., treating representation as something objective, stable across time and contexts, and quantifiable. For instance, using skin tone to measure diversity presumes skin tone is a useful proxy for race, and race is a universal category of differentiation, which may not be appropriate in all global cultures being represented. In contrast, researchers at the intersection of HCI and responsible AI have conducted qualitative evaluations of representation in genAI image models, identifying failures of representation through more participatory and community-centered methods, with evaluators who are from the same culture as the images they are evaluating (Qadri et al. 2023; Mim et al. 2024; Ghosh et al. 2024a). However, while these efforts have demonstrated the value of qualitative approaches for evaluating representation in generative image models, they do not examine what is meant by representation as a construct, nor do they critically interrogate what the participants’ process of interpretive meaning-making suggests for methods for evaluating cultural representation in AI more broadly—both of which we focus on in this paper.
@@
Recent research has shed new light on the complex process of measuring unobservable social constructs such as fairness, as well as the potential harms of reductive measurement approaches. For instance, Selbst et al. (2019) critique the definitions of fairness that lead to misleading abstractions in its evaluation. Recognizing the contested nature of the construct of fairness, Smith et al. (2024) argue for bringing in stakeholders’ expertise for more meaningful and contextually grounded evaluations. Scholars like Jacobs and Wallach (2021) advocate for utilizing approaches to measurement modeling from the social sciences to bridge the gap between unobservable constructs and their operationalization in evaluations of fairness in algorithmic systems (and for evaluation of generative AI (e.g., Wallach et al. 2025; Weidinger et al. 2025))—a framework we draw on in this paper through our discussion of the construct of cultural representation and how it is operationalized in evaluations.
@@
Jacobs and Wallach (2021) argue that measurement modeling begins with a clear and robust articulation of the construct being evaluated—in their paper, fairness, and in our paper, cultural representation. However, in prior work evaluating cultural representation in text-to-image models (e.g., Zhang et al. 2024; Kannen et al. 2024; Jha et al. 2024; Lahoti et al. 2023; Basu et al. 2023), the construct of representation is often not interrogated or only implicitly defined. Although often left implicit, prior work operationalizes representation in various ways, as lack of bias (Wan et al. 2024), as geo-cultural similarity (Basu et al. 2023), and as imbalance (Su et al. 2009). More broadly than for cultural representation in text-to-image models, Chasalow and Levy (2021) argue that researchers often fail to make explicit how they conceptualize ‘representative’ or ‘representation.’ Such various ways of operationalizing the same construct may lead to evaluations that use the same term, but optimize for disparate goals (e.g., cultural awareness, cultural diversity, avoiding “cultural bias” (Li et al. 2024), etc), making it difficult to compare the effectiveness of different approaches.
Situated Conceptualizations of Representation
@@
Decades of scholarship from the humanities have argued that representation in media, particularly visual media, cannot be defined through positivist approaches (e.g., Hall 1989, 1997; Barthes 1999). They argue that representation is not a static goal that can be objectively and quantitatively evaluated and achieved, but is instead an ongoing interpretive act whose meaning is contested and negotiated. In media studies, visual media, such as paintings, photography, or digital art, do not represent an objective reality but instead communicate multiple shades of meaning about the world (Hall 1997). That is, images have both denotative meanings (i.e., the literal subject of the image) as well as connotative meanings (i.e., the emotional subtext or symbolic meaning) the image was designed to convey—or the meaning it evokes in the viewers, which may be different than the intended meaning of the image (Hall 1997; Barthes 1999).
@@
As a result, for media studies scholars, representation in visual media is a process of meaning-making, not a one-to-one depiction of the world as it is—and thus representation is a site for struggle over meaning (Hall 1989, 1997; Desai 2000). In this way, even a photograph of an event does not objectively convey a singular true meaning (Hall 1997), but instead conveys a “positional truth” (Abu-Lughod 1991) (or multiple such truths, given the interpretive nature of images), shaped by the creators’ and viewers’ cultural and historical context. Thus, visual understanding itself relies on a slippery relationship between words, images, and the concepts signified by those images (Barthes 1999).
As media studies scholars remind us, representation in public culture is a “zone of contestation” (Desai 2000), characterized by contests over which peoples and perspectives are made visible, how they are portrayed (and by whom), contests which are shaped by who has the power to control narratives of representation. Thus, in this paper, we explore how people evaluate representation in AI-generated images and what this representation means to them in their culture.
Thick vs Thin Modes of Understanding the Social
@@
The ‘thick/thin’ dichotomy, originating in Gilbert Ryle’s philosophical work (1968), has become a cornerstone of anthropological and social scientific inquiry. He argued that understanding human behavior requires moving beyond mere observation of physical details to interpretation, recognizing the “many-layered sandwich” of meaning embedded within even seemingly simple actions (Ryle 1968). This concept was further elaborated upon by Clifford Geertz (2008), who emphasized the importance of thick description in ethnographic research. Geertz argued that researchers must not only describe actions but also interpret them within their social or cultural context. A wink, for instance, could signify a conspiratorial gesture, a flirt, a parody, or a mere rehearsal, each carrying distinct meanings depending on the social context. Sherry Ortner, in a chapter on “thick resistance,” highlighted the limitations of thin approaches in capturing the complexities of social life, which require the “richness, texture, and detail,” of thickness (Ortner 1995, 1997). She argued that thin descriptions often overlook the “internal politics of dominated groups, thin on the cultural richness of those groups, thin on the subjectivity—the intentions, desires, fears, projects—of the actors engaged in these dramas” (Ortner 1995). Ortner advocated for thick descriptions that delve into the “production, circulation, and consumption” of cultural practices, recognizing their dynamic interplay with social and political realities. More recently, Alondra Nelson has called for “thick alignment” when aligning AI systems with human values (Nelson 2023a, b). While the value of thickness in capturing the nuances of social phenomena has been widely acknowledged, scholars have also recognized the importance of thinness in discerning patterns and broader social structures. Annelise Riles Riles (2000), in her work on multi-sited ethnography, explored the challenges of achieving thick description when studying global phenomena dispersed across diverse cultures, calling attention to “the limits of thickness as a disciplinary trope” and to pay more attention to what the use of “thin composition” could yield. Anthropologist John Jackson similarly emphasized the value of thin description in providing a “baseline empiricism” and a starting point for social investigation (Jackson 2013). Thus, while thick and thin both have their place in social analysis, each one can only get you so far. In this paper, we draw on the thick/thin dichotomy to provide a framework for navigating the complexities of social inquiry when considering multiple approaches to evaluating cultural representation in AI images.
Methods